Project

General

Profile

NDOS Commissioning Meeting 2011/03/07

Attendance: Jon, Peter, Kurt, Tricia, Andrew, Chad, Art, Mat, Sue, Susan, Alena, Tom, Gavin, Rick K., Greg, Gary, Minerba, Mark Mathis, Denis

Agenda:

  1. Installation/Channel Recovery Plans
  2. DAQ Update & Plans
  3. DCS Status & Plans
  4. Shifters' report
  5. AOB

Installation/Channel Recovery:

  • Mat:
    • Installation: pressure testing of modules in block 5 and muon catcher and horizontal in block 4 will begin tomorrow.
    • Working on coming up with list of cable-ectomy. Details are in doc-db 5696. Compared this list to pixel masks; 11 of 21 that Mat picked are also heavily masked, so that might be a good place to start. Will confer with Leon.
    • all of muon catcher planes are hung with exception of "miniblock". This will be done today/tonight (overtime).

DAQ hardware:

  • Jon: would like to have a DCM ready for installation by this Thursday.
  • Rick K.: there are two DCMs that have been repaired:
    • dcm-7 was repaired and is on test-stand
    • dcm-5 was repaired and should be on test-stand
  • Ralf is coming later this week to install more PDBs

DAQ Software:

  • Gennadiy and Kurt would like to investigate the lost-trigger problem.
  • Peter: need to test functionality of new version of DDS.
    • a few minor updates related to trigger
    • last Friday, we saw empty events, and it turned out that these were all Booster spills. After some debugging, it was found that trigger validity time range was ~1 week in the past. Was related to a change in the TDU that happened around that time.
    • would be good to have trigger/monitoring scream when something like this happens.
  • To move TDU out to MINOS surf. building, need:
    • fiber optic connection to be verified (from master to slave)
    • a port allocated to us on the FNAL public network (not AD network)
    • permission to pull off accelerator signals into TDU
  • Sue: would be good to add this kind of monitoring to Ganglia?
  • Peter: interface to online monitoring/bug fixes
  • Sue: request update of DCMApplication and EventBuilderClient
  • Jon: hoping to install PostgreSQL on DAQ private dB machine
  • Kurt: request approval to work on application manager this week & snapshot service
  • Denis: updated PedestalRunner to allow multiple scans at once. Will also work on reloading kernel modules so that reboot of DCMs isn't necessary
  • Jon: kernel updates planned tomorrow morning. Peter suggested that's a good time to clean up the rats nest of cables...

DAQ Electronics:

  • Andrew: there was a plan to remap network switches and work on max. frame size. Work will commence Thursday morning.

DCS:

  • Chad placed order for RTDs, patch panels. Needs to talk to some experts re: soldering.
  • Andrew: spoke with Jeff Savage (D0 controls group) and Jim Patrick (AD controls group) re: their availability and interest. Would like to meet with experiment to discuss the current status and needs.

Any other business:

  • Chad has started working on getting MB beam information into offline
  • Tricia spoke with Laura L. about her NuSoft package for NuMI beam info... sounds like she's close to having something we can use.